A box of mass 1 kg rests on a frictionless inclined plane which is at an angle of 30◦ to the horizontal plane. Find the constant force that needs to be applied parallel to the incline to move the box down the incline with an acceleration of 1m/s$^2$.

Grade:12

1 Answers

Satyam Agarwal
5 years ago
Component of mg parallel to the incline= mgsin30 = 5 N
Let the force F acts upward parallel to the incline.
Now, 
5N -F = M*a 
F = 5- 1*1 = 4 N.